OpenStudy (anonymous):

solution set of sqrt(7x+1)-sqrt(x-5)=6

OpenStudy (anonymous):

have you done these types of problems before? it normally involves squaring both sides once, then moving some terms over, then squaring a second time.

OpenStudy (anonymous):

Ive always struggled with these solution sets

OpenStudy (anonymous):

like \[(\sqrt{7x + 1} - \sqrt{x - 5})^{2} = 36\] the left hand side becomes: \[7x + 1 - 2\sqrt{x - 5}\sqrt{7x + 1} + x - 5 = 8x - 2 \sqrt{x - 5}\sqrt{7x + 1} - 4\]

OpenStudy (anonymous):

simplified a bit becomes: \[8x - 2 \sqrt{(x - 5)(7x + 1)} - 4 = 36\]

OpenStudy (anonymous):

this was the first time we squared both sides. next we move the 8x and the -4 to the other side by subtracting 8x and adding 4: \[-2 \sqrt{(x -5)(7x + 1)} = 40 - 8x\]

OpenStudy (anonymous):

then we square both sides a second time

OpenStudy (anonymous):

\[4(x-5)(7x + 1) = (40 - 8x)^{2}\]

OpenStudy (anonymous):

expanding the right side: \[4(x - 5)(7x + 1) = 64x^{2} - 640x + 1600\] dividing everything by 4: \[(x - 5)(7x + 1) = 16x^{2} - 160x + 400\] expanding the left side: \[7x^{2} - 34x - 5 = 16x^{2} - 160x + 400\] combining like terms and rearranging: \[-9x^{2} + 126x - 405 = 0\] solve for x

OpenStudy (anonymous):

then verify the solutions you get are permitted in the domain of the original equation (because when we square, sometimes we introduce values that are not part of the original domain).

OpenStudy (anonymous):

how do you solve for x when you have to x's?

OpenStudy (anonymous):

quadratic formula works

OpenStudy (anonymous):

you could also rewrite it as \[9x^{2} - 126x + 405 = 0\] \[9x^{2} - 126x = -405\] if you dont want to use quadratic formula, you can also complete the sqaure

OpenStudy (anonymous):

5,9

OpenStudy (anonymous):

yes those are correct